Carry out the following operations as if it were a calculation of real experimental results. Express the answer with the correct number of significant figures.

 

(3.26 × 10−3 mg ) − (7.88 × 10−5 mg )


1
Expert's answer
2021-08-30T02:26:21-0400

"(326 \times 10^{-5} \;mg) -(7.88 \times 10^{-5} \;mg) \\
= 318.12 \times 10^{-5} \;mg \\
= 3.18 \times 10^{-3} \;mg"